Daniel and Lauren are painters. Daniel is faster than Lauren by 2 hours. Occasionally they work with Ian, who paints very quickly, as fast as Daniel and Lauren together. One day, they had to paint a room. Lauren and Daniel started to paint first for 1 hour, then Ian arrived an all 3 of them finished painting after another 3.5 hours. How long will it take each of them to paint the room individually?
Answer by Edwin McCravy(20056)

Daniel and Lauren are painters. Daniel is faster than Lauren
by 2 hours. Occasionally they work with Ian, who paints very
quickly, as fast as Daniel and Lauren together. One day, they
had to paint a room. Lauren and Daniel started to paint first
for 1 hour, then Ian arrived an all 3 of them finished painting
after another 3.5 hours. How long will it take each of them
to paint the room individually?
Make this chart to include all 5 situations mentioned:

               Number
              of rooms                Painting
              painted      Time         rate
             or fraction  painting       in
              thereof     in hours    rooms/hr
------------------------------------------------
Daniel         
Lauren          
Ian             
D&L          
D&L&I      

Daniel is faster than Lauren by 2 hours.
Let the time it takes Lauren to paint 1 room be L hours.
Therefore the time it takes Daniel to paint 1 room is L-2.

So we fill in 1 each for the number of rooms each can paint
and L and L-2 for their times


               Number
              of rooms                Painting
              painted      Time         rate
             or fraction  painting       in
              thereof     in hours    rooms/hr
------------------------------------------------
Daniel          1           L-2        
Lauren          1            L           
Ian             
D&L         
D&L&I

We fill in their rates in rooms/hour by dividing rooms by
hours:

               Number
              of rooms                Painting
              painted      Time         rate
             or fraction  painting       in
              thereof     in hours    rooms/hr
------------------------------------------------
Daniel          1           L-2        1/(L-2) 
Lauren          1            L           1/L
Ian             
D&L         
D&L&I   

...Ian, who paints...as fast as Daniel and Lauren together
So his rate is the sum of their rates, so we add their rates
and put that for Ian's rate and 1 for the number of rooms
he could paint at this rate.


               Number
              of rooms                Painting
              painted      Time         rate
             or fraction  painting       in
              thereof     in hours    rooms/hr
------------------------------------------------
Daniel          1           L-2        1/(L-2) 
Lauren          1            L           1/L
Ian             1                     1/(L-2)+1/L
D&L          
D&L&I

Lauren and Daniel started to paint first for 1 hour,
So we add their rates  and put that for the rate for
D&L and also put 1 hour for the time for D&L
     

               Number
              of rooms                Painting
              painted      Time         rate
             or fraction  painting       in
              thereof     in hours    rooms/hr
------------------------------------------------
Daniel          1           L-2        1/(L-2) 
Lauren          1            L           1/L
Ian             1                     1/(L-2)+1/L
D&L                          1        1/(L-2)+1/L
D&L&I     

Next we get the fraction of a room that they painted
during that 1 hour, by multiplying their rate by
their time.

then Ian arrived an all 3 of them finished painting after another 3.5 hours
So we fill in 3.5 hours for the time for D&L&I, and add the rates for
Ian and D&L.  They are the same rates so we just multiply it by 2: 


               Number
              of rooms                Painting
              painted      Time         rate
             or fraction  painting       in
              thereof     in hours    rooms/hr
------------------------------------------------
Daniel          1           L-2        1/(L-2) 
Lauren          1            L           1/L
Ian             1                     1/(L-2)+1/L
D&L          1/(L-2)+1/L     1        1/(L-2)+1/L
D&L&I                       3.5     2(1/(L-2)+1/L)

Then we fill in the fraction of the room they
painted after Ian joined in, by multiplying the
rate on the bottom row by the time:


               Number
              of rooms                Painting
              painted      Time         rate
             or fraction  painting       in
              thereof     in hours    rooms/hr
------------------------------------------------
Daniel          1           L-2        1/(L-2) 
Lauren          1            L           1/L
Ian             1                     1/(L-2)+1/L
D&L          1/(L-2)+1/L     1        1/(L-2)+1/L
D&L&I   2[3.5(1/(L-2)+1/L]  3.5     2(1/(L-2)+1/L) 

Then the equation comes from adding the fractions of
a room painted in the last two lines and setting that
equal to 1 room:

      1/(L-2)+1/L + 2[3.5(1/(L-2)+1/L)] = 1

           1/(L-2)+1/L + 7(1/(L-2)+1/L) = 1

                         8(1/(L-2)+1/L) = 1

                            8/(L-2)+8/L = 1

Multiply through by the LCD of L(L-2)

                            8L + 8(L-2) = L(L-2)
                           8L + 8L - 16 = L² - 2L
                               16L - 16 = L² - 2L    
                                      0 = L² - 18L + 16

Solve that by the quadratic formula: Get L = 17.062 hrs. 
approximately

(we also get L = 0.938 which is extraneous.)

So it takes Lauren 17.062 hrs to paint 1 room.

Since it takes 2 hours less for Daniel to paint 1 room,  

So it takes Daniel 15.062 hrs to paint 1 room.

Finally we must calculate Ian's time to paint 1 room. 

Ian's time is 1 room divided by his rate:

1 ÷ [1/(L-2)+1/L] = 
1 ÷ [1/(17.063-2)+1/17.063] =

1 ÷ 0.125

8

So it takes Ian 8 hours to paint 1 room.

Edwin